第 10 回福岡数論研究集会

本研究集会は, 一部日本学術振興会科学研究費基盤研究 (S) 16H06336, 挑戦的萌芽研究 15K13428 （研究代表者：金子昌信）の援助を受けています。

日時：2016 年 8 月 8 日 (月) − 8 月 10 日 (水) 場所：九州大学 (伊都キャンパス) ウェスト１号館

福岡市西区元岡 744

プログラム

8 月 8 日 (月) ［ＩＭＩオーディトリアム（ウエスト 1 号館 D 棟 4 階 D-413）］
10 : 15 − 11 : 05 奥村喜晶（東京工業大学）

Non-existence of certain abelian varieties with CM factors

11 : 20 − 12 : 10 小関祥康（神奈川大学）

Lattices in crystalline representations and Kisin modules associated with iterate extensions

13 : 30 − 14 : 20 島田佑一（名古屋大学）

Modularity lifting and Oda’s conjecture for Hilbert modular varieties

14 : 35 − 15 : 25 槇山賢治（京都産業大学）

A p-adic analytic family of the D-th Shintani lifting for a Coleman family and congruences between the central L-values

15 : 40 − 16 : 30 Soma Purkait （上智大学）

Minus space of half-integral weight forms

16 : 45 − 17 : 35 中屋智瑛（九州大学）

ある微分方程式のモジュラー形式解と楕円曲線の超特異多項式

8 月 9 日 (火) ［講義棟 302 （ビックリーフ 3 階, ウエスト 1 号館 B 棟 3 階と連結）］
10 : 15 − 11 : 05 川島誠（大阪大学）

G-関数の形式的 Mellin 変換であらわされる級数の特殊値の無理数性

11 : 20 − 12 : 10 中村隆（東京理科大学）

Real zeros of Hurwitz-Lerch zeta functions in the interval (−1, 0)

13 : 30 − 14 : 20 石塚裕大（京都大学）

平面曲線の行列式表示の数論的側面

14 : 35 − 15 : 25 新井啓介（東京電機大学）

ディオファントス問題と志村曲線の有理点

15 : 40 − 16 : 30 寺井伸浩（大分大学）

Euler 商に関する指数型不定方程式について

16 : 45 − 17 : 35 小松尚夫（Wuhan University）

Euler numbers and related numbers, and their applications

19 : 00 −  懇親会
8 月 10 日 (水) ［講義棟 302 （ビックリーフ 3 階, ウエスト 1 号館 B 棟 3 階と連結）］
10 : 15 − 11 : 05 廣津孝（東北大学）

Brauer groups of Châtelet surfaces over p-adic fields

11 : 20 − 12 : 10 Yu Yang （京都大学）

Local p-rank and semi-stable reduction of curves

13 : 30 − 14 : 20 澤田晃一郎（京都大学）

Pro-p and cohomological aspects of anabelian geometry of hyperbolic polycurves

14 : 35 − 15 : 25 南出新（京都大学）

配置空間の遠アーベル幾何について

※ 研究集会の最新情報は下記アドレスでご確認ください。

  http://auemath.aichi-edu.ac.jp/~ykishi/FSNT/16-fukuokaNT.html

世話人：金子昌信（九州大学) mkaneko[at]math.kyushu-u.ac.jp
権寧魯（九州大学) ygon[at]math.kyushu-u.ac.jp
岸康弘（愛知教育大学)ykishi[at]auecc.aichi-edu.ac.jp